OpenStudy (anonymous):

If a single card is drawn from a standard 52-card deck, what is the probability that it is neither a King nor a club?

terenzreignz (terenzreignz):

Well, it might be more prudent to first check the probability that it IS a king or a club. How many cards in the deck are either Kings or clubs?

OpenStudy (anonymous):

I believe in a deck there are 4 kings and 13 clubs including the king of clubs

terenzreignz (terenzreignz):

So, that makes how many? 17 ?

OpenStudy (anonymous):

i think 16 because u dont include the one club that is a king

terenzreignz (terenzreignz):

I was testing @kkbrant , @allea_marie but ok :P So what's the probability that you pick a club or a King, out of a standard deck of 52 cards?

OpenStudy (anonymous):

I think it'll be 16/52 but you reduce it to 4/13 ?

OpenStudy (anonymous):

16 out of 52 however the question is that the card we draw is not a king or a club I got the same thing at @allea_marie but you would remove them out of the deck 52-4 kings-12 clubs=35 so do you think it would 35/52???

OpenStudy (anonymous):

36/52

terenzreignz (terenzreignz):

No :) Here's the rule, given an event x, and p(x) being the probability that x happens the probability of x NOT happening is 1-p(x)

terenzreignz (terenzreignz):

That is to say, to get the probability that something DOESN'T happen, get the probability that it DOES happen, then subtract it from 1.

OpenStudy (anonymous):

I dont know lol...i think it is 35/52 because 52-1=51 then subtract 13 because there is 13 cards per suit and subtract 3 because there are 4 kinds but one of which are a club...so you get 35...So 35 of the 52 are neither king nor club.

OpenStudy (mertsj):

There are 16 unsuccessful outcomes and 36 successful outcomes. So the probability that it is a club or a king is 16/52. The probability that it is NOT a club or a king is 36/52

terenzreignz (terenzreignz):

We can do this the easy way or the hard way, guys (girls) Let's look at hearts... there are 13 of them, but one of them is a king, so only 12 are 'allowed' Same goes for diamonds, 12 of them are 'allowed' as there is one king of diamonds Same goes for spades, 12 of them are 'allowed', as there is one king of spades. None of the clubs are allowed. So that makes 12+12+12 allowable draws. =36. Out of 52 cards. I do not recommend doing it like that :D

OpenStudy (anonymous):

@terenzreignz you confuse me lol

terenzreignz (terenzreignz):

Exactly. That way is confusing. What I basically did was literally count all the cards that are neither kings nor clubs.

OpenStudy (anonymous):

So i can see

terenzreignz (terenzreignz):

So, you now agree that it's 36/52 = 9/13 ?

OpenStudy (mertsj):

And what happened to the asker is this discussion?

terenzreignz (terenzreignz):

Must have understood already? Always assume the best of people :)

OpenStudy (anonymous):

im still trying to understand, I keep getting 35/52 I apologize its just not making sense yet @Mertsj what am i missing @terenzreignz that I am only getting 35?? How are you getting 36?

OpenStudy (anonymous):

@kkbrant im with you...Im still getting 35

OpenStudy (mertsj):

There are 13 clubs and 3 kings that are not clubs. That makes 16 cards that are clubs and/or kings. Do you understand that?

OpenStudy (mertsj):